# AT 2025abao: the fourth luminous red nova in M 31

### v2

@@ 3.2.2 LRN outburst
In the case of AT 2025abao, this phase is followed by a fast brightening of about five magnitudes in the -band (six in ): this is the onset of the LRN outburst. The rapid rise lasts about one week from the discovery of AT 2025abao. Then, its light curve settles on a sort of plateau phase during which the object maintains a nearly constant brightness ( mag). To constrain the timing of the main LRN maximum, we perform a 2nd-order polynomial fit to the -band data collected between MJD = 60973 and MJD = 61000, determining that the light curve peaked on MJD 60986.30.2, at a magnitude of mag. This epoch will be used as a reference for all phases considered in this paper. We note that the bluer-band light curves peak earlier ( d in , d in ), while a photometric peak is not well-defined in the red bands, where a plateau is noticeable, lasting for about two months. We measured the post-maximum decline rates in the optical bands by fitting the light curve in each filter with a straight line. The decline is steepest in the -band ( mag 100 d-1), and less and less steep as we move towards redder filters (, , mag 100 d-1). The -band light curve instead is flat, and the -band one is actually still rising, up to phase d, when it reached its maximum. The plateau phase ends at d and is followed by a sudden fall of the light curve, and then by a linear decline.
Individual patterns of LRN light curves are not fully understood. However, there is a general consensus that the characteristic double-peaked shape is due to a merging event (Banerjee et al., 2026). As a consequence of the common envelope ejection and the loss of systemic angular momentum, the two components of the binary system may merge. If this happens, the main LRN outburst is produced by the violent ejection of material following the coalescence of the two stellar cores, and its subsequent interaction with the pre-existing common envelope (see, e.g., Metzger and Pejcha, 2017; MacLeod et al., 2017, 2022; Matsumoto and Metzger, 2022; Kirilov et al., 2025).
@@ 3.3 Colour evolution
The intrinsic colours of AT 2025abao, blue soon after its discovery ( mag at d), evolve to redder colours after the -band maximum ( mag at d, mag at d). Thanks to the scanning by the ZTF survey we have, for the first time, an indication of the optical colour of a LRN during the pre-outburst slow luminosity rise. The colour stays roughly around 1.4 mag one year before the -band maximum. Then, it reaches a maximum up to mag (with large scatter) at the time of the -band minimum ( months). decreases to 1.1 mag at d, at the start of the fast rise to the first peak, diminishing to a lowest of mag at d. From there, increases again to 1.3 mag at d, at the time of the plateau drop. The colour curves of AT 2025abao are shown in Figure 3.
@@ 3.4 Comparisons with similar objects
As comparison objects, we select a few faint LRNe well monitored during the slow pre-outburst rise phase. Given that AT 2025abao has an extensive and well-sampled light curve since a few years before the LRN event (K25), our primary reference object is V1309 Sco (Mason et al., 2010; Tylenda et al., 2011), although the two objects have significantly different light curves during the outburst. AT 2025abao reaches an absolute magnitude at the -band peak of mag (Fig. 4), hence it lies at the faint edge of the extragalactic LRNe brightness distribution, although it is significantly more luminous than Galactic objects such as V1309 Sco and V4332 Sgr (Pastorello et al., 2019). AT 2019zhd (in M 31; Pastorello et al., 2021a) and the Galactic V838 Mon (Goranskij et al., 2020, and references therein) have similar absolute magnitude at maximum as AT 2025abao, and are hence selected as comparison objects.
@@
As mentioned in Sect. 3.2.1, the pre-outburst light curves of both AT 2025abao and V1309 Sco were very well monitored (see Fig. 4). While the former was much brighter at the top of the slowly rising phase (about one year before the LRN outburst onset), reaching mag, the latter reached only mag. Then, both objects dimmed towards the optically thick common envelope phase. In particular, AT 2025abao reached a minimum at mag at d, before starting the fast rise towards the LRN maximum light, when it increased its brightness by about 6 mag.
@@
From the comparison in Fig. 4, we note that the fast rise of AT 2025abao to the main peak is very similar to that exhibited by AT 2019zhd, in both speed and luminosity. At peak, the absolute magnitude of AT 2025abao in the band is only marginally brighter than those of AT 2019zhd ( mag) and V838 Mon ( mag; note that the Figure shows the -band light curve, which peaks at mag). Then, its light curve shows a flat, long-lasting (about two months) plateau, while the comparison objects show a short-duration (less than a week) early peak before declining.V838 Mon shows a more structured post-maximum light curve, with a later rebrightening (e.g., Munari et al., 2002). In contrast to the comparison objects, the outburst of V1309 Sco is much fainter, and a shift upwards by 2.5 mag has to be applied to its -band absolute light curve in Fig. 4 to match the peak luminosity of AT 2025abao.
@@
Due to the plateau-like feature after the first peak, the light curve of AT 2025abao is reminiscent of other LRNe that do not exhibit an evident second red peak, instead showing a long phase of nearly constant luminosity. Known LRNe with plateau-like light curves are AT 2021afy (Pastorello et al., 2023), although it is 4.5 mag more luminous and at the bright edge of the LRNe brightness distribution; AT 2018bwo, whose plateau reaches mag (Blagorodnova et al., 2021; Pastorello et al., 2023); AT 2020hat, that peaks at mag (Pastorello et al., 2021b). However, we remark that AT 2018bwo was discovered already after the first peak due to the Solar conjunction, hence the object could have reached a higher peak luminosity.
@@ 3.5 Bolometric light curve
Given that only some photometric bands are available before the discovery and the start of the follow-up campaign of the main outburst, we divided the dataset in two parts and constructed two different bolometric light curves: (1) We first constructed a pseudo-bolometric curve selecting only the bands, integrating their fluxes (inferred by converting the magnitudes taking into account for the distance and reddening) with the trapezoidal rule from to days, hence covering the pre-discovery phases. (2) Then, we built the bolometric curve of AT 2025abao extended from the UV to the NIR, accounting for the contribution from to bands, starting at the discovery epoch, hence covering only the main LRN outburst. For epochs without observations in some bands, we interpolated the available data using the -band light curve as a reference and assuming a constant colour index. Both the pseudo- and the fully bolometric curves of AT 2025abao are shown in Fig. 5.
The bolometric luminosity of AT 2025abao at the time of the -band maximum is , similar to the values reached by the comparison objects. However, its top luminosity () is reached about 67 d later, just before the fall from the plateau. This peak bolometric luminosity makes AT 2025abao the brightest LRN ever discovered in M 31 and, at the time of the peak, one of the most luminous objects within that galaxy. However, while the bolometric luminosity of AT 2025abao at peak is similar to that of the comparison objects and the fast rise resembles that of AT 2019zhd, the later evolution is completely different, as none of the LRNe considered presents a 70 d-lasting plateau. In this respect, AT 2025abao is more similar to AT 2018bwo (Blagorodnova et al., 2021; Pastorello et al., 2023), although the former is about three times fainter. The total radiated energy, calculated by integrating the UV+optical+NIR bolometric curve between phases and d, is erg. This value has to be considered as a lower limit since we are missing the contributions from wavelengths longer than 2.5 m, and the energy emitted prior to phase d, i.e. the fast rise before the peak of the outburst. If we add the contribution from the filters from to between phases and d, and from the filters between and d, we obtain a total radiated energy of erg, which still has to be taken as an underestimation.
@@ 3.6 SED evolution
We constructed the Spectral Energy Distribution (SED) of AT 2025abao at different epochs, starting from phase d, combining the photometric data from the NUV to the NIR domains. Then, we fitted them with a Planckian function to estimate how the black-body (BB) temperature () and the radius () evolve with time. Between d and d, we did so using only the data from the filters from to . For epochs without the photometric information in some filter, we made interpolations or extrapolations from adjacent epochs with available data, assuming no colour changes. The highest temperature was reached on d (the epoch of the minimum colour), at K. At the time of the -band maximum light (see Fig. 6), the SED peaks at around 5000 Å and is fitted with a single BB function having a temperature of K. The inferred photospheric radius is . At these early phases, no IR excess is detected. In Fig. 7, we plot the temporal evolution of , , and the BB luminosity . During the plateau phase, the temperature slowly decreases and flattens at 4300 K, slightly lower than that expected for the H recombination, as evidenced also by the redder shape of the spectral continuum in this phase (see Sect. 4). The photospheric radius when the object was hottest is cm (corresponding to ), while during the plateau phase it slowly expands to cm (hence, ). During the plateau phase, slightly increases from at the time of the -band maximum to erg s-1 about 70 d later, before a sharp decrease after the fall from the plateau.
3.7 SPHEREx SED of the precursor
@@
The Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx; Bock et al., 2025) is a space telescope launched on 11 March 2025 to conduct an all-sky survey in 102 infrared colours between 0.75 and 5 m, hence covering a wide range of wavelengths from the NIR to the mid infrared (MIR) domains. During its first scanning of the whole celestial sphere, between MJD 60865 and 60902 ( d to d), it observed the field of AT 2025abao during the pre-LRN phase, at a time close to the optical minimum. We retrieved the calibrated SED (Akeson et al., 2025; Hui et al., 2026), from the Spectrophotometry Tool available in the NASA/IPAC Infrared Science Archive webpage121212https://irsa.ipac.caltech.edu/applications/spherex/tool-spectrophotometry. The IR SED of the precursor of AT 2025abao from SPHEREx is presented in Figure 8.
@@
This provides a rare opportunity to analyse a NIR/MIR spectrum of a precursor of an LRN, and probably for the first time, the MIR region is observed in spectroscopy. The source is faint in the bluest bands and shows a red continuum between 0.74 and 1.2 m. Three peaks are clearly present at 1.3, 1.7, and 2.2 m. These peaks, and the broad absorptions in between due to molecular absorption bands from TiO, H2O, and CO, are also visible in the NIR spectra presented by K25 and taken in 2024 (see their Figure 3). Then, at m, the SED is nearly flat in the MIR, which could be due to thermal emission from dust surrounding the progenitor system of AT 2025abao. For reference, in Fig. 8 we plot together a NIR/MIR spectrum of a M7e III star (HD 108849, a known AGB131313https://simbad.u-strasbg.fr/simbad/sim-basic?Ident=bk+vir) from the Atlas by Rayner et al. (2009), which perfectly matches the SED from SPHEREx. A modelling of the SED is presented in Sect. 5. Finally, we performed spectro-photometry on the SED to derive approximate NIR and MIR magnitudes (in , , , , WISE and bands) of AT 2025abao 3-4 months before the -band maximum. The infrared colours are larger, hence redder, than during the outburst (for instance, mag vs. mag at d).
3.8 SPHEREx SED of the outburst
@@
Between MJD 61027 and 61041 ( d to d), SPHEREx observed again the field of AT 2025abao, this time during the outburst stage. The corresponding IR SED, presented in Figure 9, shows an hot thermal continuum, with absorption bands at 0.82 m (visible also in the optical spectra, produced by VO and TiO molecules, as observed also in V838 Mon; Lynch et al. 2004), 0.94 m (due to TiO; Valenti et al., 1998), 1.1 m (from a combination of VO, CN and TiO molecules; Pastorello et al., 2023), 1.45 m (water vapours), and 2.4 m (from CO; Blagorodnova et al., 2021). We performed a Planckian fit to the reddening-corrected SED, which provides an excellent match up to the red tail with a BB having the following parameters: K, , and erg s-1. These parameters are consistent with those obtained from the optical filters at similar phases. No IR excess is detected up to 5 m, as a single BB is sufficient to fit the data even at the reddest wavelengths. Therefore, either dust is not present and had yet to be formed, or it is pre-existing and is much colder than 600 K, otherwise its emission would have been visible as an IR excess at such long wavelengths.
4 Spectral evolution
@@
AT 2025abao was classified as a LRN on 2025 October 29, 10 days after the official discovery date and at a phase of d (Taguchi and Maeda, 2025b). However, earlier spectra from amateur astronomers are also available. The earliest, taken on 2025 October 27 (phase d), was presented by Skopal et al. (2025). One day later, another low-resolution spectrum was obtained by C. Balcon (phase d), and additional spectra were obtained with a nearly daily cadence (at phases , and d) by Taguchi and Maeda (2025b) and Skopal et al. (2025). These early spectra exhibit a blue continuum ( K) with superimposed narrow emission lines of the Balmer series, a feature typical of LRNe during the rise to maximum light (Pastorello et al., 2019). The classification spectrum of Taguchi and Maeda (2025b), retrieved from the TNS, is shown in Fig. 10, where it is compared with spectra of LRNe AT 2019zhd and V1309 Sco at the maximum light to emphasize their overall similarity. Nonetheless, the spectral evolution of V1309 Sco is more rapid, and metal lines are already visible in absorption at these early epochs.
@@
We conducted a spectroscopic follow-up of AT 2025abao, during which we collected 21 optical spectra spanning more than three months of evolution. The log of spectroscopic observations is given in Table 2, and the spectral time series is shown in Figure 11, left panel. The spectra from the 1.22m Galileo telescope equipped with a Boller & Chivens (B&C) spectrograph and the 1.82m Copernico telescope with an Echelle spectrograph were reduced with a custom-made, IRAF-based pipeline developed by A. Siviero, while the spectra from the 2.56m Nordic Optical Telescope (NOT) with ALFOSC, the 1.82m Copernico telescope with AFOSC, and the 10.4m Gran Telescopio Canarias (GTC) plus OSIRIS were reduced using the foscgui141414Foscgui is a graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. pipeline. Finally, the spectra of the 3.58m Telescopio Nazionale Galileo (TNG) plus DOLORES spectra were hand-reduced using routine IRAF procedures.

### v3

@@ 3.2.2 LRN outburst
In the case of AT 2025abao, this phase is followed by a fast brightening of about five magnitudes in the band (6 in ); this is the onset of the LRN outburst. The rapid rise lasts about one week from the discovery of AT 2025abao. Then, its light curve settles on a sort of plateau phase during which the object maintains a nearly constant brightness ( mag). To constrain the timing of the main LRN maximum, we performed a second-order polynomial fit to the -band data collected between MJD = 60973 and MJD = 61000, determining that the light curve peaked on MJD 60986.30.2, at a magnitude of mag. We used this epoch as a reference for all phases considered in this paper. We note that the bluer band light curves peak earlier ( d in , d in ), while a photometric peak is not well-defined in the red bands, where a plateau is noticeable, lasting for about two months. We measured the post-maximum decline rates in the optical bands by fitting the light curve in each filter with a straight line. The decline is steepest in the band ( mag 100 d-1), and less and less steep as we move towards redder filters (, , mag 100 d-1). The -band light curve instead is flat, and the -band one is actually still rising, up to phase d, when it reached its maximum. The plateau phase ends at d; it is followed by a sudden fall of the light curve and, 20 days later, by a linear decline with a slope of mag 100 d-1.
Individual patterns of LRN light curves are not fully understood. However, there is a general consensus that the characteristic double-peaked shape is due to a merging event (Banerjee et al., 2026). As a consequence of the common envelope ejection and the loss of systemic angular momentum, the two components of the binary system may merge. If this happens, the main LRN outburst is produced by the violent ejection of material following the coalescence of the two stellar cores, and its subsequent interaction with the pre-existing common envelope (e.g. Metzger and Pejcha, 2017; MacLeod et al., 2017, 2022; Matsumoto and Metzger, 2022; Kirilov et al., 2025; Chen, 2026).
@@ 3.3 Colour evolution
The intrinsic colours of AT 2025abao, which are blue soon after its discovery ( mag at d), evolve to redder colours after the -band maximum ( mag at d, mag at d). Thanks to the scanning by the ZTF survey we have, for the first time, an indication of the optical colour of an LRN during the pre-outburst slow luminosity rise. The colour stays roughly around 1.4 mag one year before the -band maximum. Then, it reaches a maximum up to mag (with large scatter) at the time of the -band minimum ( months). decreases to 1.1 mag at d, at the start of the fast rise to the first peak, diminishing to a lowest of mag at d. From there, increases again to 1.3 mag at d, at the time of the plateau drop. The colour curves of AT 2025abao are shown in Fig. 3.
@@ 3.4 Comparisons with similar objects
As comparison objects, we select a few faint LRNe well monitored during the slow pre-outburst rise phase. Given that AT 2025abao has an extensive and well-sampled light curve since a few years before the LRN event (K25), our primary reference object is V1309 Sco (Mason et al., 2010; Tylenda et al., 2011), although the two objects have significantly different light curves during the outburst. AT 2025abao reaches an absolute magnitude at the -band peak of mag (Fig. 4), hence it lies at the faint edge of the extragalactic LRNe brightness distribution, although it is significantly more luminous than Galactic objects such as V1309 Sco and V4332 Sgr (Pastorello et al., 2019). AT 2019zhd (in M 31; Pastorello et al., 2021a) and the Galactic V838 Mon (Goranskij et al. 2020 and references therein) have similar absolute magnitudes to AT 2025abao at their maximum, and they were hence selected as comparison objects.
@@
As mentioned in Sect. 3.2.1, the pre-outburst light curves of both AT 2025abao and V1309 Sco were very well monitored (see Fig. 4). While the former was much brighter at the top of the slowly rising phase (about one year before the LRN outburst onset), reaching mag, the latter reached only mag. Then, both objects dimmed towards the optically thick common-envelope phase. In particular, AT 2025abao reached a minimum at mag at d, before starting the fast rise towards the LRN maximum light, when it increased its brightness by about 6 mag.
@@
From the comparison in Fig. 4, we note that the fast rise of AT 2025abao to the main peak is very similar to that exhibited by AT 2019zhd in terms of both speed and luminosity. At the peak, the absolute magnitude of AT 2025abao in the band is only marginally brighter than those of AT 2019zhd ( mag) and V838 Mon ( mag; note that the figure shows the -band light curve, which peaks at mag). Then, its light curve shows a flat, long-lasting (about two months) plateau, while the comparison objects show a short-duration (less than a week) early peak before declining. V838 Mon shows a more structured post-maximum light curve, with a later re-brightening (e.g. Munari et al., 2002). In contrast to the comparison objects, the outburst of V1309 Sco is much fainter, and a shift upwards by 2.5 mag has to be applied to its -band absolute light curve in Fig. 4 to match the peak luminosity of AT 2025abao.
@@
Due to the plateau-like feature after the first peak, the light curve of AT 2025abao is reminiscent of other LRNe that do not exhibit an evident second red peak, instead showing a long phase of nearly constant luminosity. Known LRNe with plateau-like light curves are AT 2021afy (Pastorello et al., 2023), although it is 4.5 mag more luminous and at the bright edge of the LRNe brightness distribution; AT 2018bwo, whose plateau reaches mag (Blagorodnova et al., 2021; Pastorello et al., 2023); and AT 2020hat, which peaks at mag (Pastorello et al., 2021b). However, we remark that AT 2018bwo was already discovered after the first peak due to the solar conjunction; hence, the object could have reached a higher peak luminosity.
@@ 3.5 Bolometric light curve
Given that only some photometric bands are available before the discovery and the start of the follow-up campaign of the main outburst, we divided the dataset into two parts and constructed two different bolometric light curves. We first constructed a pseudo-bolometric curve selecting only the and bands, integrating their fluxes (inferred by converting the magnitudes taking into account for the distance and reddening) with the trapezoidal rule from to days, hence covering the pre-discovery phases. Then, we built the bolometric curve of AT 2025abao extended from the UV to the NIR, accounting for the contribution from to the bands, starting at the discovery epoch, hence only covering the main LRN outburst. For epochs without observations in some bands, we interpolated the available data using the -band light curve as a reference and assuming a constant colour index. Both the pseudo- and the fully bolometric curves of AT 2025abao are shown in Fig. 5.
The bolometric luminosity of AT 2025abao at the time of the -band maximum is , similar to the values reached by the comparison objects. However, its top luminosity () was reached about 67 d later, just before the fall from the plateau. This peak bolometric luminosity makes AT 2025abao the brightest LRN ever discovered in M 31 and, at the time of the peak, one of the most luminous objects within that galaxy. However, while the bolometric luminosity of AT 2025abao at the peak is similar to that of the comparison objects and the fast rise resembles that of AT 2019zhd, the later evolution is completely different, as none of the LRNe considered present a plateau lasting 70 days. In this respect, AT 2025abao is more similar to AT 2018bwo (Blagorodnova et al., 2021; Pastorello et al., 2023), although the former is about three times fainter. The total radiated energy, calculated by integrating the UV+optical+NIR bolometric curve between phases and d, is erg. This value has to be considered as a lower limit since we are missing the contribution emitted at wavelengths longer than 2.5 m and shorter than 2000 Å, and prior to phase d, i.e. at the time of the fast rise before the maximum light. If we add the contribution of to bands between phases and d, and of the filters between and d, we obtain a total radiated energy of erg.
@@ 3.6 SED evolution
We constructed the spectral energy distribution (SED) of AT 2025abao at different epochs, starting from phase d, combining the photometric data from the NUV to the NIR domains. Then, we fitted them with a Planckian function to estimate how the black-body (BB) temperature () and the radius () evolve with time. Between d and d, we did so using only the data from the filters from to . For epochs without the photometric information in a given filter, we made interpolations or extrapolations from adjacent epochs with available data, assuming no colour changes. The highest temperature was reached on d (the epoch of the minimum colour), at K. At the time of the -band maximum light (see Fig. 6), the SED peaks at around 5000 Å and is fitted with a single BB function that has a temperature of K. The inferred photospheric radius is . At these early phases, no IR excess is detected. In Fig. 7, we plot the temporal evolution of , , and the BB luminosity . During the plateau phase, the temperature slowly decreases and flattens at 4300 K, slightly lower than expected for the H recombination, as also evidenced by the redder shape of the spectral continuum in this phase (see Sect. 4). The photospheric radius when the object was hottest is cm (corresponding to ), while during the plateau phase it slowly expands to cm (hence, ). During the plateau phase, slightly increases from at the time of the -band maximum to erg s-1 about 70 d later, before a sharp decrease occurs after the fall from the plateau.
3.7 SPHEREx observations
3.7.1 SED of the precursor
@@
The Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx; Bock et al., 2026) is a space telescope launched on 2025 March 11 to conduct an all-sky survey in 102 infrared colours between 0.75 and 5 m, hence covering a wide range of wavelengths from the NIR to the mid-infrared (MIR) domains. During its first scanning of the whole celestial sphere, between MJD 60865 and 60902 ( d to d), it observed the field of AT 2025abao during the pre-LRN phase, at a time close to the optical minimum. We retrieved the public calibrated SED (Akeson et al., 2025; Hui et al., 2026) from the spectrophotometry tool available on the NASA/IPAC Infrared Science Archive web page.121212https://irsa.ipac.caltech.edu/applications/spherex/tool-spectrophotometry The IR SED of the precursor of AT 2025abao from SPHEREx is presented in Fig. 8.
@@
This provides a rare opportunity to analyse an NIR+MIR spectrum of a precursor of an LRN, and the MIR region is observed in spectroscopy for the first time. The source is faint in the bluest bands and shows a red continuum between 0.74 and 1.2 m. Three peaks are clearly present at 1.3, 1.7, and 2.2 m. These peaks, and the broad absorptions in between due to molecular absorption bands from TiO, H2O, and CO, are also visible in the NIR spectra presented by K25 and taken in 2024 (see their Figure 3). Then, at m, the SED is nearly flat in the MIR, which could be due to thermal emission from dust surrounding the progenitor system of AT 2025abao. For reference, in Fig. 8 we plot an NIR+MIR spectrum of a M7e III star (HD 108849, a known AGB131313https://simbad.u-strasbg.fr/simbad/sim-basic?Ident=bk+vir) from the Atlas by Rayner et al. (2009), which perfectly matches the SED from SPHEREx. A modelling of the SED is presented in Sect. 5. Finally, we performed spectro-photometry on the SED to derive approximate NIR and MIR magnitudes (in , , , , WISE and bands) of AT 2025abao 3–4 months before the -band maximum. The infrared colours are larger (thus redder) than during the outburst (i.e. mag vs. mag at d).
3.7.2 SED of the outburst
@@
Between MJD 61027 and 61041 ( d to d), SPHEREx observed the field of AT 2025abao again, this time during the outburst. The corresponding IR SED, presented in Fig. 9, shows a hot thermal continuum, with absorption bands at 0.82 m (analogous to those seen in the optical spectra, produced by VO and TiO molecules, as also observed in V838 Mon; Lynch et al. 2004), 0.94 m (due to TiO; Valenti et al., 1998), 1.1 m (from a combination of VO, CN, and TiO molecules; Pastorello et al., 2023), 1.45 m (water vapours) and 2.4 m (from CO; Blagorodnova et al., 2021). We performed a Planckian fit to the reddening-corrected SED, which provides an excellent match up to the red tail with a BB that has the following parameters: K, , and erg s-1. These parameters are consistent with those obtained from the optical filters at similar phases. No IR excess is detected up to 5 m, as a single BB is sufficient to fit the data even at the reddest wavelengths. Therefore, either dust is not present and has yet to form, or there is pre-existing dust, much colder than 600 K, otherwise its emission would have been visible as an IR excess at such long wavelengths.
4 Spectroscopic evolution
@@
The transient AT 2025abao was classified as an LRN on 2025 October 29, ten days after the official discovery date and at a phase of d (Taguchi and Maeda, 2025b). However, earlier spectra from amateur astronomers are also available. The earliest, taken on 2025 October 27 (phase d), was presented by Skopal et al. (2025). One day later, another low-resolution spectrum was obtained by C. Balcon (phase d), and additional spectra were obtained with a nearly daily cadence (at phases , , and d) by Taguchi and Maeda (2025b) and Skopal et al. (2025). These early spectra exhibit a blue continuum ( K) with superimposed narrow emission lines of the Balmer series, a feature typical of LRNe during the rise to maximum light (Pastorello et al., 2019). The classification spectrum of Taguchi and Maeda (2025b), retrieved from the TNS, is shown in Fig. 10, where it is compared with spectra of LRNe AT 2019zhd and V1309 Sco at the maximum light to emphasise their overall similarity. Nonetheless, the spectral evolution of V1309 Sco is more rapid, and metal lines are already visible in absorption at these early epochs.
@@
We conducted a spectroscopic follow-up of AT 2025abao, during which we collected 21 optical spectra spanning more than three months of evolution. The log of spectroscopic observations is given in Table 3, and the spectral time series is shown in Fig. 11, left panel. The spectra from the 1.22m Galileo telescope equipped with a Boller & Chivens (B&C) spectrograph and the 1.82m Copernico telescope with an Echelle spectrograph were reduced with a custom-made, IRAF-based pipeline developed by A. Siviero, while the spectra from the 2.56m Nordic Optical Telescope (NOT) with ALFOSC, the 1.82m Copernico telescope with AFOSC, and the 10.4m Gran Telescopio Canarias (GTC) plus OSIRIS were reduced using the foscgui141414Foscgui is a graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. pipeline. Finally, the spectra of the 3.58m Telescopio Nazionale Galileo (TNG) plus DOLORES spectra were hand-reduced using routine IRAF procedures.
4.1 Main spectral evolution